OpenStudy (anonymous):

What is the value of x, when 10(x + 2) = 5(x + 8)? A) x = 1 B) x = 2 C) x = 4 D) x = 5/6

OpenStudy (denonakavro):

10(x+2)=5(x+8) 10x+20=5x+40 5x=20 X=4

OpenStudy (er.mohd.amir):

open braket

OpenStudy (anonymous):

Lets solve: Step 1: Simplify both sides of the equation. 10(x+2)=5(x+8) Simplify: (Hide steps) (10)(x)+(10)(2)=(5)(x)+(5)(8)(Distribute) 10x+20=5x+40 Step 2: Subtract 5x from both sides. 10x+20−5x=5x+40−5x 5x+20=40 Step 3: Subtract 20 from both sides. 5x+20−20=40−20 5x=20 Step 4: Divide both sides by 5. \[\frac{5x }{ 5 }=\frac{ 20 }{ 5 }\] Answer: x=4

OpenStudy (anonymous):

thanks yall